OpenStudy (anonymous):

Ball 1, with a mass of 10kg, is moving to the right at 20 m/s. Ball 2 is moving to the left at 20m/s. Upon collision, ball 2 comes to a standstill, and ball 1 moves to the left at twice its original speed. What is the mass of ball 2? The mass of ball 2 is ?????…..

OpenStudy (jamesj):

Use Conservation of Momentum. Write p1 for the momentum of the first ball, p2 for the second; write p1' and p2' for the momentum afterwards. The by Conservation of Momentum: p1 + p2 = p1' + p2' We know that p2' = 0, because p2' = m2.v2' and v2' = 0 because the ball is still. We also know that p1 = m1.v1 = 10(20) = 200 and p2 = m2.v2 = 20.m2 Further, because the first ball goes off at twice its original speed, v2' = 40 m/s and hence p2' = 10(40) = 400 Hence, bringing all of that together and using Conservation of Momentum, p1 + p2 = p1' + p2' 200 + 20.m2 = 400 + 0 Now solve for m2.